# E-TRIALS: Empowering Data-Driven Decisions to Enhance Computer-Based Learning Platforms

### v1

@@ 5 Results & Discussions
In this section, we present the results of the example experiments. Tables 3, 3, and 4 summarize the ATE estimates, $\hat{\tau}_{t-test}$, $\hat{\tau}_{OLS}$, and $\hat{\tau}_{LOOP}$, along with their respective variation and significance measures .
@@
First, we present the results of the analysis for Experiment 14, shown as Table 3, which aims to determine whether providing students with feedback that either aligns closely with the problem at hand or diverges from its context influences their proximal and distal-term performance. We include the estimates, standard errors, and $p_{value}$ of the results, accompanied by a figure visualizing the confidence intervals. This visualization helps us explore whether there is any likely significant positive or negative effect.
Table 3: Experiment 14: Similar vs. Different Example Feedback
Performance Method Estimate Std. Error $\mathbf{Pr(>|t|)}$
Proximal $\hat{\tau}_{t-test}$ 0.0254 0.0333 0.4461
$\hat{\tau}_{OLS}$ 0.0233 0.0290 0.4240
$\hat{\tau}_{LOOP}$ 0.0073 0.0318 0.8201
Distal $\hat{\tau}_{t-test}$0.0053 0.0285 0.8516
$\hat{\tau}_{OLS}$0.0141 0.0199 0.4814
$\hat{\tau}_{LOOP}$0.0219 0.0237 0.3584
  • Note: The smallest Std.Error in each performance group is in bold.
Figure 6: Experiment 14: Confidence Intervals
@@
From Table 3, we observe no significant difference between the similar and different feedback groups. However, after covariate adjustment, the standard errors, as indicated by the Std.Error column, are reduced. The results show that $\hat{\tau}_{OLS}$ achieves the smallest standard error for both proximal and distal performance. It is important to note, though, that $\hat{\tau}_{OLS}$ is biased, whereas $\hat{\tau}_{LOOP}$ is unbiased and exhibits the second smallest standard error. These findings confirm that while both OLS regression and LOOP provide more precise estimators compared to student’s t-test, only LOOP offers an unbiased estimate. Thus, it is reasonable to select LOOP as the preferred estimator for estimating ATE. From Figure 6, we observe that all confidence intervals include $y=0$, indicating that none of the estimates are statistically significant. However, there is a noticeable trend: for proximal performance, students receiving different feedback tend to perform better, whereas for distal performance, they tend to perform worse. This is reflected in the proportion of the spread of the confidence intervals, which leans positive for proximal performance and negative for distal performance. Additionally, we can see that the lengths of the confidence intervals for $\hat{\tau}_{OLS}$ and $\hat{\tau}_{LOOP}$ are shorter than those for the $\hat{\tau}_{t-test}$, suggesting that OLS and LOOP provide more precise estimates compared to the student’s t-test.
Next, we interpret the results of Experiment 15, which examines the impact of different visualization techniques: drawing, mental visualization, and provided images on students’ performance in both proximal and distal terms. The calculated results are presented in Table 4, with corresponding visualizations shown in Figures 7, 8, and 9.
Table 4: Experiment 15: Drawing vs. Imaging vs. Image Provided
Condition Performance Method Estimate Std. Error $\mathbf{Pr(>|t|)}$
Drawing vs. Imaging Proximal $\hat{\tau}_{t-test}$0.0465 0.0352 0.1897
$\hat{\tau}_{OLS}$0.0469 0.0291 0.1091
$\hat{\tau}_{LOOP}$0.0469 0.0291 0.1091
Distal $\hat{\tau}_{t-test}$0.0293 0.0323 0.3651
$\hat{\tau}_{OLS}$0.0352 0.0198 0.0774
$\hat{\tau}_{LOOP}$0.0352 0.0198 0.0774
Imaging vs. Image Provided Proximal $\hat{\tau}_{t-test}$ 0.0472 0.0360 0.1919
$\hat{\tau}_{OLS}$ 0.0377 0.0309 0.2238
$\hat{\tau}_{LOOP}$ 0.0377 0.0309 0.2238
Distal $\hat{\tau}_{t-test}$ 0.0559 0.0330 0.0928
$\hat{\tau}_{OLS}$ 0.0405 0.0189 0.0346
$\hat{\tau}_{LOOP}$ 0.0405 0.0189 0.0346
Drawing vs. Image Provided Proximal $\hat{\tau}_{t-test}$ 0.0008 0.0331 0.9814
$\hat{\tau}_{OLS}$0.0025 0.0297 0.9341
$\hat{\tau}_{LOOP}$0.0025 0.0297 0.9341
Distal $\hat{\tau}_{t-test}$ 0.0266 0.0288 0.3575
$\hat{\tau}_{OLS}$ 0.0196 0.0177 0.2719
$\hat{\tau}_{LOOP}$ 0.0196 0.0177 0.2719
  • Note: The smallest standard error within each performance group and the significant $p$-values are in bold.
Figure 7: Experiment 15: Confidence Intervals
Figure 8: Experiment 15: Confidence Intervals
Figure 9: Experiment 15: Confidence Intervals
In Table 4, we observe that $\hat{\tau}_{OLS}$ and $\hat{\tau}_{LOOP}$ provide the same estimates and standard errors to four decimal places. Both methods exhibit smaller standard errors compared to $\hat{\tau}_{t-test}$. However, it is important to note that $\hat{\tau}_{LOOP}$ is unbiased, making it the recommended estimator. An interesting finding in this experiment is that, when comparing the Imaging vs. Image Provided groups, both the OLS regression and LOOP methods produce significant $p_{value}<0.05$. However, after applying a multiple pairwise comparison adjustment, the significance of these $p_{value}$ disappears. Nonetheless, this result suggests that the two groups may exhibit a meaningful difference if we can further improve the precision of our ATE estimators.
We examine Figures 7, 8, and 9. The results indicate that Imaging tends to negatively impact students’ performance compared to Drawing in both proximal and distal periods. In contrast, Image Provided appears to enhance students’ performance in both periods. Additionally, while Drawing shows no significant difference compared to Image Provided in proximal performance, Image Provided demonstrates a positive effect on students’ distal performance compared to Drawing.
@@
Finally, we examine the insights gained from Experiment 16, which investigated whether prompting students to reflect on their confusion before accessing a hint has impact on students’ performance. The findings are summarized in Table 5 and illustrated in Figure 10.
Table 5: Experiment 16: Reflect vs. Regular Hint
Performance Method Estimate Std. Error $\mathbf{Pr(>|t|)}$
Proximal $\hat{\tau}_{t-test}$ 0.0055 0.0248 0.8237
$\hat{\tau}_{OLS}$ 0.0072 0.0174 0.6776
$\hat{\tau}_{LOOP}$ 0.0101 0.0171 0.5543
Distal $\hat{\tau}_{t-test}$0.0060 0.0191 0.7555
$\hat{\tau}_{OLS}$0.0083 0.0109 0.4435
$\hat{\tau}_{LOOP}$0.0069 0.0113 0.5428
  • Note: The smallest standard error within each performance group is in bold.
Figure 10: Experiment 16: Confidence Intervals
@@
Similar to the findings from the previous two experiments, $\hat{\tau}_{LOOP}$ is the recommended estimator due to its combination of a small standard error and unbiased property. Additionally, Figure 10 suggests that prompting students to reflect appears to enhance their performance during the proximal period but negatively impacts their performance in the distal term.
@@
By reviewing these three experiments, we observe the clear advantages of using $\hat{\tau}_{LOOP}$ to estimate ATE. This estimator is both precise and easy to apply, making it a versatile choice for other experiments involving comparative tutoring designs. One notable observation is that the distal performance estimates have smaller standard errors and narrower confidence intervals. At first glance, this might seem counterintuitive, as distal performance represents a long-term imputation. However, this phenomenon makes sense when we consider that proximal performance was included as a super covariate in the estimation of distal performance. Since students’ current performance is highly predictive of their future performance, this inclusion improves the precision of the distal performance estimates.
@@ 6.1 Limitations
In this paper, we utilized data from E-Trials to evaluate whether different tutoring strategies have a significant positive or negative impact on students’ proximal and distal performance, where LOOP estimator are recommended to estimate ATE due to its reduced standard error. However, there are certain limitations associated with the methods employed.
@@
    The ATE estimation approach used in this paper was specifically tailored for support comparison experiments conducted in E-TRIALS. It won’t work for single support or problem varied studies, or studies from a different CBLP.
@@
    For the LOOP estimator to achieve a small standard error, it is crucial that $\hat{m}_{i}$ serves as a reliable estimator of $m_{i}$. Ensuring their closeness poses a significant challenge that must be addressed.
@@
    The LOOP method cannot handle categorical variables with more than $53$ levels, as observed when including $class\_xid$ as a covariate to adjust for students’ baseline performance. Further revisions to the LOOP package are necessary to overcome this limitation.
@@ 6.2 Future Plans
The method and code discussed in this paper are specifically designed for support comparison studies. One future direction is to develop a similar framework for single support studies. However, for problem-varied studies, creating a generalized method is less feasible due to the flexibility these studies provide researchers in designing and conducting experiments. This autonomy makes it challenging to standardize the process, as each study may require unique adaptations.
Another future plan is to build a Shiny application to fully automate the data cleaning and modeling process. This application would allow users to upload their CSV files and select, from a predefined list, their desired outcome variable. By automating this process, the Shiny application could reduce time and effort required to conduct experiments.
@@ 6.3 Conclusion
In this paper, we introduce E-TRIALS, a tool designed to help researchers make data-driven decisions to improve student performance in CBLPs. E-TRIALS supports multiple study types, each with its own advantages and disadvantages, enabling researchers to conduct experiments tailored to their specific needs. Using data from three support comparison studies, we developed a generalizable approach that can be applied to any support comparison experiment within E-TRIALS. Our work demonstrates the potential of E-TRIALS to advance research in the realm of learning sciences and contribute to the development of more effective, inclusive, and adaptive CBLPs.

### v2

@@ 5 Results & Discussions
In this section, we present the results of the examples of experiments. Tables 5(a) and 5(b) summarize the ATE estimates, $\hat{\tau}_{\text{t-test}}$, $\hat{\tau}_{\text{Reg}}$, and $\hat{\tau}_{\text{LOOP}}$, together with their respective variations and significance measures.
@@
First, we present the results of the analysis for Experiment 1, shown in Table 5(a), which aims to determine whether providing students with feedback that aligns closely with the problem at hand or diverges from its context influences their proximal and distal performance. We include estimates, standard errors, and $p-value$ of the results, accompanied by a Figure 4 visualizing the confidence intervals. This visualization helps us explore whether there is a likely significant positive or negative effect.
Table 5: Experiment 1 & 2: Comparison of Treatment Effect Estimation
(a) Experiment 1: Similar vs. Different Feedback
Perf. Method Est. SE p-value
Prox. $\hat{\tau}_{t\text{-test}}$ 0.0254 0.0333 0.4461
$\hat{\tau}_{\text{Reg}}$ 0.0207 0.0290 0.4757
$\hat{\tau}_{\text{LOOP}}$ 0.0042 0.0316 0.8957
Distal $\hat{\tau}_{t\text{-test}}$0.0053 0.0285 0.8516
$\hat{\tau}_{\text{Reg}}$0.0187 0.0198 0.3443
$\hat{\tau}_{\text{LOOP}}$0.0270 0.0233 0.2474
Smallest SE per group is in bold.
(b) Experiment 2: Reflect vs. Regular Hint
Perf. Method Est. SE p-value
Prox. $\hat{\tau}_{t\text{-test}}$ 0.0055 0.0248 0.8237
$\hat{\tau}_{\text{Reg}}$ 0.0084 0.0174 0.6297
$\hat{\tau}_{\text{LOOP}}$0.0196 0.0173 0.2594
Distal $\hat{\tau}_{t\text{-test}}$0.0060 0.0191 0.7555
$\hat{\tau}_{\text{Reg}}$0.0073 0.0108 0.5031
$\hat{\tau}_{\text{LOOP}}$0.0136 0.0115 0.2373
Smallest SE per group is in bold.
Figure 4: Experiment 1: Confidence Intervals
@@
From Table 5(a), we observe no significant difference between the similar and different feedback groups. However, after covariate adjustment, the standard errors, as indicated by the Std.Error column, are reduced. The results show that $\hat{\tau}_{\text{Reg}}$ achieves the smallest standard error for both proximal and distal performance. However, it is important to note that $\hat{\tau}_{\text{Reg}}$ might be biased, while $\hat{\tau}_{\text{LOOP}}$ is unbiased and exhibits the second smallest standard error. These findings confirm that while both Regression and LOOP provide more precise estimators compared to the student’s t-test, only LOOP offers an unbiased estimate based on equation 10. Thus, it is reasonable to select LOOP as the preferred estimator to estimate ATE. From Figure 4, we observe that all confidence intervals include $y=0$, indicating that none of the estimates are statistically significant. However, there is a noticeable trend: for proximal performance, students receiving different feedback tend to perform better, whereas for distal performance, they tend to perform worse. This is reflected in the proportion of the spread of the confidence intervals, which leans positive for proximal performance and negative for distal performance. Additionally, we can see that the lengths of the confidence intervals for $\hat{\tau}_{Reg}$ and $\hat{\tau}_{LOOP}$ are shorter than those for the $\hat{\tau}_{t-test}$, suggesting that regression and LOOP provide more precise estimates compared to the student’s t-test.
@@
Next, we examine the insights gained from Experiment 2, which investigated whether prompting students to reflect on their confusion before accessing a hint has an impact on students’ performance. The findings are summarized in Table 5(b) illustrated in Figure 5.
Figure 5: Experiment 2: Confidence Intervals
@@
Similar to the findings from the previous experiment, $\hat{\tau}_{LOOP}$ is the recommended estimator due to its combination of a small standard error and unbiased property. Additionally, Figure LABEL:fig:exp1416_comparison right hand suggests that prompting students to reflect is more likely to enhance their performance in the proximal period, but may negatively affect their performance in the distal term—although this effect is not statistically significant at the 5% level.
@@
By reviewing these two experiments, we observe the clear advantages of using $\hat{\tau}_{LOOP}$ to estimate ATE. This estimator is both precise and easy to apply, making it a versatile choice for other experiments involving comparative tutoring designs. One notable observation is that the distal performance estimates have smaller standard errors and narrower confidence intervals. At first glance, this may seem counterintuitive, given that distal performance involves long-term prediction, which is typically more challenging. However, this phenomenon becomes understandable when considering that proximal performance—a variable highly correlated with distal outcomes—was included as a super covariate in the distal performance model. Since students’ current performance strongly predicts their future performance, incorporating this information enhances the precision of the distal performance estimates.
@@ 6.1 Limitations
In this paper, we utilized data from E-TRIALS to evaluate whether different tutoring strategies have a significant positive or negative impact on students’ proximal and distal performance, where the LOOP estimator is recommended for estimating the ATE due to its reduced standard error and unbiasedness. However, there are certain limitations associated with the methods employed.
@@
    Although the vanilla ATE estimation code works well for support-comparison RCTs with student-level randomization, modifications are required when randomization is applied at a higher level (e.g., class or school). Additional adjustments may also be necessary for single-support studies, problem-varied studies, or experiments conducted in a different CBLP. However, as long as the provided data follow the E-TRIALS format, these methods can still be applied effectively.
@@
    For the LOOP estimator to achieve low variance, it is essential that $\hat{m}_{i}$ is a reliable estimate of $m_{i}$, which requires accurate estimation of both $c_{i}$ and $t_{i}$. Additionally, to ensure a variance reduction relative to $\hat{\tau}_{t-test}$, it is necessary that $m_{i}\perp T_{i}$. Satisfying both conditions can sometimes be challenging.
@@
    The LOOP method cannot handle categorical variables with more than $53$ levels, as observed when including $class\_xid$ as a covariate to adjust for the baseline performance of the students in Experiment 2. Further revisions to the LOOP package are necessary to overcome this limitation.
@@ 6.2 Future Plans
The method and code discussed in this paper are designed to support comparison studies. One future direction is to develop a similar framework for single support studies. However, it is not feasible to create a generalized framework for problem-varied studies due to the flexibility these studies provide researchers in designing and conducting experiments. This autonomy makes it challenging to standardize the process, as each study may require unique adaptations.
Another future plan is to build a Shiny application to fully automate the data cleaning and modeling process. This application would allow users to upload their CSV files and select from a predefined list of outcome variables. By automating this process, the Shiny application could reduce the time and effort required to evaluate experiments.
@@ 6.3 Conclusion
In this paper, we introduced E-TRIALS, a tool designed to help researchers make data-driven decisions to improve student performance in CBLPs. E-TRIALS supports multiple study types, each with its own advantages and disadvantages, enabling researchers to conduct experiments tailored to their specific needs. Our work demonstrates the potential of E-TRIALS to advance research in the realm of learning sciences and support the design of evidence-based CBLPs.